# Euclid Quick Data Release (Q1)

### v1

@@ 2.1 Q1 photometry and spectroscopy
The Q1 (Q1-TP001; Q1cite) data set contains images and photometric catalogues from both VIS ($I_{\scriptscriptstyle\rm E}$ band) and NISP ($Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ bands), and one-dimensional spectra of the NISP spectroscopic mode.
In this work, we use the main photometric catalogue (catalogue.mer_catalogue in the Euclid Science Archive [^1] https://eas.esac.esa.int/sas/) generated by the MERge Processing Function (MER; Q1-TP004), which include aperture fluxes, template-fit and Srsic-fit fluxes, and quality flags in each band, as well as morphological information for all sources detected in the Euclid Deep Fields. We also use the main morphology catalogue (catalogue.mer_morphology; Q1-SP040) to assess the morphological properties of the final sample.
Combined one-dimensional (1D) slitless spectroscopic data of sources brighter than $H_{\scriptscriptstyle\rm E}=22.5$ have been generated by the dedicated SIR spectroscopic processing function (Q1-TP006). The slitless spectra used in this work are retrieved from ESA Datalabs [^2] https://datalabs.esa.int/ (Datalabscite).
Because the observed slitless spectrum of an object is the convolution of its intrinsic spectrum with its wavelengthdependent spatial light profile along the dispersion direction, sources with larger extents will have lower spectral resolution. In addition, the blue and red ends of the observed spectrum tend to show upturns due to the convolution nature of the observed spectrum, and the lower signal-to-noise at both ends (see e.g., 2004ApJS..154..501P). Each of the original 1D spectra contains 531 data points, covering 11 900–19 002 Å with a wavelength interval of 13.4 Å. To keep only the useful data from the slitless spectra, we trim the blue and red ends, retaining 12 047–18 734 Å, which yields 500 data points per spectrum.
@@ 2.2 Reliable quasar candidates from Gaia and AllWISE
Gaia DR3 announced a sample of 6.6 million quasar candidates (the qso_candidates table, hereafter the GDR3 QSO candidate catalogue; GaiaCollaboration_2023_2023A&A...674A...1G; GaiaCollaboration_2023_2023A&A...674A..41G), which has high completeness thanks to the combination of several different modules, including the Discrete Source Classifier (DSC), the Quasar Classifier (QSOC), the variability classification module, the surface brightness profile module, and the Gaia DR3 Celestial Reference Frame source table. The DSC uses the Gaia Blue (BP) and Red (RP) Photometer (2023A&A...674A...2D) spectrum together with the mean $G$-band magnitude, the variability in this band, the parallax, and the proper motion to classify each Gaia source probabilistically into five classes: quasars, galaxies, stars, white dwarfs, and physical binary stars. The QSOC determines the redshifts using BP/RP spectra of the sources classified as quasars by the DSC (2023A&A...674A..31D). The variability classification module identifies 25 classes of variable sources (including AGN candidates) from the variability of the Gaia light curves using supervised machine learning (2023A&A...674A..14R; 2023A&A...674A..24C). Both the surface-brightness profile module and the Gaia DR3 Celestial Reference Frame source table are based on external catalogues of quasars and quasar candidates (see 2023A&A...674A..11D; 2022A&A...667A.148G, for a complete list). Despite its high completeness, the GDR3 QSO candidate catalogue has an estimated low purity of quasars (52 %) and a large scatter of redshift estimates (GaiaCollaboration_2023_2023A&A...674A..41G).
@@
    Quaia (Storey-Fisher2024) with nearly 1.3 million sources at $G<20.5$. This sample is selected using a set of cuts involving proper motion, Gaia and UnWISE (2019ApJS..240...30S) colours and magnitudes, which are designed to remove stellar contaminants of the Milky Way and the Large and Small Magellanic Clouds.
@@
    CatNorth (Fu2024_CatNorth) with more than 1.5 million sources down to the Gaia limiting magnitude ($G<21.0$) in the 3 $\pi$ sky of the Pan-STARRS1 (PS1; Chambers2016) footprint ($\delta>-30\degr$). This catalogue is primarily built with a machine learning classification model trained on colour and morphological features from Gaia, PS1, and CatWISE2020 (2021ApJS..253....8M). An additional probabilistic cut on proper motion (probability density of zero proper motion; see 2021ApJS..254....6F; Fu2024_CatNorth, for the definition) is applied to further purify the candidates. The CatNorth catalogue has a purity of approximately 90 %.
@@
    CatSouth (Fu2025_catsouth) with 0.9 million sources with $G<21.0$ covered by the fourth data release (DR4) of the SkyMapper Southern Survey (SMSS; $\delta\lesssim 16\degr$; Onken_2024_2024PASA...41...61O). This catalogue is built with the same method as CatNorth, while based on data from Gaia, SMSS DR4, and VISTA (Visible and Infrared Survey Telescope for Astronomy) surveys (2006Msngr.126...41E; 2010NewA...15..433M; 2011A&A...527A.116C; 2013Msngr.154...35M; 2013Msngr.154...32E), as well as CatWISE2020.
All three purified catalogues propagate the Gaia DR3 QSOC template-matching redshift and combine it with multi-band photometry to derive improved photometric redshifts. These photometric redshifts reduce the fractions of catastrophic outliers by more than 15 % compared to the original QSOC redshift, at the cost of a modest decrease in precision (Storey-Fisher2024; Fu2024_CatNorth). The compilation of the three catalogues above contains more than 1.9 million unique (with unique Gaia source_id) quasar candidates in the entire sky. This purified GDR3 QSO candidate catalogue will be referred to as GDR3-QSOs hereafter for simplicity. Together, 4467 sources are matched to the Q1 main photometric catalogue (see Sect. 2.1 for details) using Gaia source_id, which is listed in both GDR3-QSOs and catalogue.mer_catalogue of Q1.
@@
Table 1: Summary of numbers of input sources and identified quasars. The success rate is the number of identified quasars divided by that of SIR sources.
@@ 3.1 Template-matching redshifts for the candidates
For each observed quasar spectrum, we first estimate the redshift by comparing it to a rest-frame template via a Pearson correlation function (PCF, also known as the normalised cross-correlation) method (2018A&A...616A...6S). We begin with the template from Glikman_2006, based on 27 bright quasars observed with the NASA Infrared Telescope Facility (IRTF). After constructing a new composite spectrum in Sect. 4.3, we re-run the PCF with the new composite as a template to refine the redshift estimates; for the final measurements we adopt a piecewise template that combines 2001AJ....122..549V, the mean composite from this work, and Glikman_2006, see Sect. B.2. The PCF algorithm proceeds as follows:
@@
    The PCF is computed over the set $\mathcal{I}$ of overlapping wavelength bins using the median-normalised spectra above:
@@
    To account for the logarithmic nature of wavelength shifts, the redshift grid is sampled with a step size that scales with $(1+z)$:
@@
    where $\delta$ is a base step (e.g., $\delta=0.001$ in this work).
@@ 3.2 Visual inspection
An interactive visual inspection tool PGSpecPlot, part of the Python package specbox [^3] https://github.com/rudolffu/specbox, has been used to check the spectra of the quasar candidates. During visual inspection, each spectrum is displayed sequentially with an overplotted quasar template adjusted to the estimated template redshift ($z_{\mathrm{temp}}$). Using interactive controls, including a slider with non-linear $(1+z)$ scaling and a corresponding spin box (a text box with an up-down control), the user can verify the template redshift and adjust it as necessary. Keyboard shortcuts facilitate the rapid classification of each spectrum (e.g., flagging non-quasar objects or uncertain cases). A history of inspected spectra is maintained in a comma-separated value file so that previously processed spectra can be automatically loaded.
@@
In addition to the initial $z_{\mathrm{temp}}$, a Gaia redshift is displayed in the window when available. The Gaia redshift (hereafter $z_{\textit{Gaia}}$) is primarily based on the Gaia DR3 low-resolution spectral template-matching redshift from QSOC (redshift_qsoc; 2023A&A...674A..31D; GaiaCollaboration_2023_2023A&A...674A..41G), supplemented by the photometric redshift from CatNorth and CatSouth when redshift_qsoc is not available. Because $z_{\textit{Gaia}}$ and $z_{\mathrm{temp}}$ are independent spectroscopic redshift estimates obtained from different wavelength ranges ($z_{\textit{Gaia}}$ from the BP/RP bands within 330–1050 nm, and $z_{\mathrm{temp}}$ from ${\rm RG}_{\scriptscriptstyle\rm E}$ in the NIR), a $z_{\mathrm{temp}}$ that is close to $z_{\textit{Gaia}}$ (with $|z_{\mathrm{temp}}-z_{\textit{Gaia}}|/[1+z_{\mathrm{temp}}]<0.15$) is taken as a secure visual redshift ($z_{\mathrm{vi}}$) in most cases, even when only one emission line is present in the wavelength range of ${\rm RG}_{\scriptscriptstyle\rm E}$. Nevertheless, when $z_{\textit{Gaia}}$ is unavailable and only one emission line is seen, the redshift can be highly uncertain. Such single-line spectra are labelled as ‘uncertain-redshift’ objects unless the emission line and continuum fit the template with high confidence during visual inspection.
@@
As expected from the predictions (EP-Lusso), quasars with $0.89<z<1.83$ are most easily identified with the H $\alpha$ emission line, which is the strongest and broadest line among all emission lines detected by ${\rm RG}_{\scriptscriptstyle\rm E}$. Quasars at lower redshift ($z<0.89$) are mainly identified with the combination of several rest-frame NIR emission lines, i.e., [S iii] $\lambda$9071, Pa $\delta$, He i + Pa $\gamma$ (blended), and Pa $\beta$. At $1.83<z<2.85$, quasars are identified with H $\beta$, [O iii], and H $\gamma$. At $2.85<z<3.3$, the quasar spectra lack strong features except for H $\gamma$ and the pseudo-continuum (‘small blue bump’), and they are mainly identified with the agreement between $z_{\mathrm{temp}}$ and $z_{\textit{Gaia}}$. At $z>3.3$, Mg ii enters the wavelength range, and the redshift determination is secured with the combination of $z_{\mathrm{temp}}$ (from the Mg ii line) and $z_{\textit{Gaia}}$. Figure 1 shows example Euclid spectra of identified quasars at different rest-frame wavelength ranges, which correspond to the different emission line features described above.
@@
Approximately 2300 sources are labelled as ‘uncertain-redshift’ objects, making up 25 % of the entire sample (9214 spectra). Given the complexity of the data reduction and limited depth of the slitless spectroscopy, there are also approximately 3400 unidentifiable spectra with: (i) corrupted data with too many invalid or anomalously high flux values; or (ii) featureless spectra either due to weak lines or low S/N. These spectra are currently classified as unknown sources. However, given the high reliability of our input catalogues, it is likely that many of these objects are indeed quasars that could be confirmed with deeper observations or improved data processing in the future.
@@
3.3 Spurious redshift rejection and the final redshift
@@
To evaluate the performance of our visual redshift $z_{\mathrm{vi}}$, we first examine the source concentration parameter $\mu_{\mathrm{max}}-\mathrm{mag}$ (mumax_minus_mag in catalogue.mer_catalogue). Here, $\mu_{\mathrm{max}}$ is the source peak surface brightness above the background, and $\mathrm{mag}$ is the magnitude used to compute point-like probability, both given by SourceXtractor++ [^4] https://github.com/astrorama/SourceXtractorPlusPlus (2020ASPC..527..461B; 2022arXiv221202428K) during the MER data reduction. The estimator $\mu_{\mathrm{max}}-\mathrm{mag}$ is related to the concentration of light at the peak versus the total magnitude; at a given magnitude, sources with smaller $\mu_{\mathrm{max}}-\mathrm{mag}$ are more point-like (Q1-TP004). This parameter has also been used in 2012MNRAS.426.3369J, 2022arXiv220705709S, 2023A&A...671A.146E, Q1-SP027, and Q1-SP003 as input for point/extended source classification.
@@
Figure 2 shows the distribution of visually identified quasars in the $\mu_{\mathrm{max}}-\mathrm{mag}$-$z_{\mathrm{vi}}$ plane. In general, $\mu_{\mathrm{max}}-\mathrm{mag}$ decreases as $z_{\mathrm{vi}}$ becomes higher. Most sources exhibit compact morphologies (low $\mu_{\mathrm{max}}-\mathrm{mag}$ values) at $z_{\mathrm{vi}}\gtrsim 0.8$, consistent with unresolved point-like sources. However, a distinct subset of sources at $z_{\mathrm{vi}}>2$ shows significantly higher $\mu_{\mathrm{max}}-\mathrm{mag}$ values, indicative of extended or low-concentration profiles that are atypical for high-redshift quasars. These sources are flagged as spurious redshifts and excluded from the final sample using a conservative empirical cut: $\mu_{\mathrm{max}}-\mathrm{mag}>-2$ at $z_{\mathrm{vi}}>2$, as indicated by the red dashed lines. This selection removes sources whose morphology and redshift are inconsistent with expectations for high-redshift quasars, improving the redshift accuracy of the final catalogue.
@@
Figure 2: Concentration parameter $\mu_{\mathrm{max}}-\mathrm{mag}$ as a function of the visual redshift $z_{\mathrm{vi}}$. The rejected region is shaded in pink, and is defined with $\mu_{\mathrm{max}}-\mathrm{mag}>-2$ at $z_{\mathrm{vi}}>2$, as indicated by the red dashed lines. Sources inside the rejected region are considered to have spurious redshifts and are marked with black crosses.

### v2

@@ 2.1 Q1 photometry and spectroscopy
The Q1 (Euclid Collaboration: Aussel et al. 2025; Euclid Quick Release Q1 2025) dataset contains images and photometric catalogues from both VIS ($I_{\scriptscriptstyle\rm E}$ band) and NISP ($Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ bands), and one-dimensional (1D) spectra of the NISP spectroscopic mode. In this work, we use the main photometric catalogue (catalogue.mer_catalogue in the Euclid Science Archive [^1] https://eas.esac.esa.int/sas/) generated by the MERge Processing Function (MER; Euclid Collaboration: Romelli et al. 2025), which include aperture fluxes, template-fit and Sérsic-fit fluxes, and quality flags in each band, as well as morphological information for all sources detected in the Euclid Deep Fields. We also use the main morphology catalogue (catalogue.mer_morphology; Euclid Collaboration: Quilley et al. 2025) to assess the morphological properties of the final sample.
Combined 1D slitless spectroscopic data of sources brighter than $H_{\scriptscriptstyle\rm E}=22.5$ have been generated by the dedicated SIR spectroscopic processing function (Euclid Collaboration: Copin et al. 2025). The slitless spectra used in this work are retrieved from ESA Datalabs [^2] https://datalabs.esa.int/ (Navarro et al. 2024).
Because the observed slitless spectrum of an object is the convolution of its intrinsic spectrum with its wavelength‑dependent spatial light profile along the dispersion direction, sources with larger extents will have lower spectral resolution. In addition, the blue and red ends of the observed spectrum tend to show upturns due to the convolution nature of the observed spectrum, and the lower S/N at both ends (see e.g. Pirzkal et al. 2004). Each of the original 1D spectra contains 531 data points, covering 11 900–19 002 Å with a wavelength interval of 13.4 Å. To keep only the useful data from the slitless spectra, we trim the blue and red ends, retaining 12 047–18 734 Å, which yields 500 data points per spectrum.
@@ 2.2 Reliable quasar candidates from Gaia and AllWISE
Gaia DR3 announced a sample of 6.6 million quasar candidates (the qso_candidates table, hereafter the GDR3 QSO candidate catalogue; Gaia Collaboration: Vallenari et al. 2023; Gaia Collaboration: Bailer-Jones et al. 2023), which has high completeness thanks to the combination of several different modules, including the Discrete Source Classifier (DSC), the Quasar Classifier (QSOC), the variability classification module, the surface brightness profile module, and the Gaia DR3 Celestial Reference Frame source table. The DSC uses the Gaia Blue (BP) and Red (RP) Photometer (De Angeli et al. 2023) spectrum together with the mean $G$-band magnitude, the variability in this band, the parallax, and the proper motion to classify each Gaia source probabilistically into five classes: quasars, galaxies, stars, white dwarfs, and physical binary stars. The QSOC determines the redshifts using BP and RP spectra of the sources classified as quasars by the DSC (Delchambre et al. 2023). The variability classification module identifies 25 classes of variable sources (including AGN candidates) from the variability of the Gaia light curves using supervised machine learning (Rimoldini et al. 2023; Carnerero et al. 2023). Both the surface-brightness profile module and the Gaia DR3 Celestial Reference Frame source table are based on external catalogues of quasars and quasar candidates (see Ducourant et al. 2023; Gaia Collaboration: Klioner et al. 2022, for a complete list). Despite its high completeness, the GDR3 QSO candidate catalogue has an estimated low purity of quasars (52 %) and a large scatter of redshift estimates (Gaia Collaboration: Bailer-Jones et al. 2023).
@@
    Quaia (Storey-Fisher et al. 2024) with nearly 1.3 million sources at $G<20.5$. This sample is selected using a set of cuts involving proper motion, Gaia and UnWISE (Schlafly et al. 2019) colours and magnitudes, which are designed to remove stellar contaminants of the Milky Way and the Large and Small Magellanic Clouds.
@@
    CatNorth (Fu et al. 2024) with more than 1.5 million sources down to the Gaia limiting magnitude ($G<21.0$) in the 3 $\pi$ sky of the Pan-STARRS1 (PS1; Chambers et al. 2016) footprint ($\delta>-30\degr$). This catalogue is primarily built with a machine learning classification model trained on colour and morphological features from Gaia, PS1, and CatWISE2020 (Marocco et al. 2021). An additional probabilistic cut on proper motion (probability density of zero proper motion; see Fu et al. 2021, 2024, for the definition) is applied to further purify the candidates. The CatNorth catalogue has a purity of approximately 90 %.
@@
    CatSouth (Fu et al. 2025) with 0.9 million sources with $G<21.0$ covered by the fourth data release (DR4) of the SkyMapper Southern Survey (SMSS; $\delta\lesssim 16\degr$; Onken et al. 2024). This catalogue is built with the same method as CatNorth, while based on data from Gaia, SMSS DR4, and VISTA (Visible and Infrared Survey Telescope for Astronomy) surveys (Emerson et al. 2006; Minniti et al. 2010; Cioni et al. 2011; McMahon et al. 2013; Edge et al. 2013), as well as CatWISE2020.
All three purified catalogues propagate the Gaia DR3 QSOC template-matching redshift and combine it with multi-band photometry to derive improved photometric redshifts. These photometric redshifts reduce the fractions of catastrophic outliers by more than 15 % compared to the original QSOC redshift, at the cost of a modest decrease in precision (Storey-Fisher et al. 2024; Fu et al. 2024). The compilation of the three catalogues above contains more than 1.9 million unique (with unique Gaia source_id) quasar candidates in the entire sky. This purified GDR3 QSO candidate catalogue will be referred to as GDR3-QSOs hereafter for simplicity. Together, 4467 sources are matched to the Q1 main photometric catalogue (see Sect. 2.1 for details) using Gaia source_id, which is listed in both GDR3-QSOs and catalogue.mer_catalogue of Q1.
@@
Table 1: Summary of numbers of input sources and identified quasars. [^3] The success rate is the number of identified quasars divided by that of SIR sources.
@@ 3.1 Template-matching redshifts for the candidates
For each observed quasar spectrum, we first estimated the redshift by comparing it to a rest-frame template via a Pearson correlation function (PCF, also known as the normalised cross-correlation) method (Sartoretti et al. 2018). We began with the template from Glikman et al. (2006), based on 27 bright quasars observed with the NASA Infrared Telescope Facility (IRTF). After constructing a new composite spectrum in Sect. 4.3, we re-ran the PCF with the new composite as a template to refine the redshift estimates; for the final measurements we adopted a piecewise template that combines Vanden Berk et al. (2001), the mean composite from this work, and Glikman et al. (2006), see Sect. C.2. The PCF algorithm proceeds as follows:
@@
    The PCF is computed over the set, $\mathcal{I}$, of overlapping wavelength bins using the median-normalised spectra above:
@@
    To account for the logarithmic nature of wavelength shifts, the redshift grid is sampled with a step size that scales with $1+z$:
@@
    where $\delta$ is a base step (e.g. $\delta=0.001$ in this work).
@@ 3.2 Visual inspection
An interactive visual inspection tool PGSpecPlot, part of the Python package specbox (Fu 2026), was used to check the spectra of the quasar candidates. During visual inspection, each spectrum is displayed sequentially with an overplotted quasar template adjusted to the estimated template redshift ($z_{\mathrm{temp}}$). Using interactive controls, including a slider with non-linear $(1+z)$ scaling and a corresponding spin box (a text box with an up-down control), the user can verify the template redshift and adjust it as necessary. Keyboard shortcuts facilitate the rapid classification of each spectrum (e.g. flagging non-quasar objects or uncertain cases). A history of inspected spectra is maintained in a comma-separated value file so that previously processed spectra can be automatically loaded.
@@
In addition to the initial $z_{\mathrm{temp}}$, a Gaia redshift is displayed in the window when available. The Gaia redshift (hereafter $z_{\textit{Gaia}}$) is primarily based on the Gaia DR3 low-resolution spectral template-matching redshift from QSOC (redshift_qsoc; Delchambre et al. 2023; Gaia Collaboration: Bailer-Jones et al. 2023), supplemented by the photometric redshift from CatNorth and CatSouth when redshift_qsoc is not available. Because $z_{\textit{Gaia}}$ and $z_{\mathrm{temp}}$ are independent spectroscopic redshift estimates obtained from different wavelength ranges ($z_{\textit{Gaia}}$ from the BP or RP bands within 330–1050 nm, and $z_{\mathrm{temp}}$ from ${\rm RG}_{\scriptscriptstyle\rm E}$ in the NIR), a $z_{\mathrm{temp}}$ that is close to $z_{\textit{Gaia}}$ (with $|z_{\mathrm{temp}}-z_{\textit{Gaia}}|/[1+z_{\mathrm{temp}}]<0.15$) is taken as a secure visual redshift ($z_{\mathrm{vi}}$) in most cases, even when only one emission line is present in the wavelength range of ${\rm RG}_{\scriptscriptstyle\rm E}$. Nevertheless, when $z_{\textit{Gaia}}$ is unavailable and only one emission line is seen, the redshift can be highly uncertain. Such single-line spectra are labelled as ‘uncertain-redshift’ objects unless the emission line and continuum fitted the template with high confidence during visual inspection.
@@
As expected from the predictions (Euclid Collaboration: Lusso et al. 2024), quasars with $0.89<z<1.83$ are most easily identified with the H $\alpha$ emission line, which is the strongest and broadest line among all emission lines detected by ${\rm RG}_{\scriptscriptstyle\rm E}$. Quasars at lower redshift ($z<0.89$) are mainly identified with the combination of several rest-frame NIR emission lines, i.e. [S iii] $\lambda$9071, Pa $\delta$, He i + Pa $\gamma$ (blended), and Pa $\beta$. At $1.83<z<2.85$, quasars are identified with H $\beta$, [O iii], and H $\gamma$. At $2.85<z<3.3$, the quasar spectra lack strong features except for H $\gamma$ and the pseudo-continuum (‘small blue bump’), and they are mainly identified with the agreement between $z_{\mathrm{temp}}$ and $z_{\textit{Gaia}}$. At $z>3.3$, Mg ii enters the wavelength range, and the redshift determination is secured with the combination of $z_{\mathrm{temp}}$ (from the Mg ii line) and $z_{\textit{Gaia}}$. Figure 1 shows example Euclid spectra of identified quasars at different rest-frame wavelength ranges, which correspond to the different emission line features described above.
@@
Approximately 2300 sources are labelled as ‘uncertain-redshift’ objects, making up 25 % of the entire sample (9214 spectra). Given the complexity of the data reduction and limited depth of the slitless spectroscopy, there are also approximately 3400 unidentifiable spectra with: (i) corrupted data with too many invalid or anomalously high flux values; or (ii) featureless spectra due to either weak lines or low S/N. These spectra are currently classified as unknown sources. However, given the high reliability of our input catalogues, it is likely that many of these objects are indeed quasars that could be confirmed with deeper observations or improved data processing in the future.
@@
3.3 Spurious redshift rejection, consistency check with DESI, and the final redshift
@@
To evaluate the performance of our visual redshift $z_{\mathrm{vi}}$, we first examined the source concentration parameter $\mu_{\mathrm{max}}-\mathrm{mag}$ (mumax_minus_mag in catalogue.mer_catalogue). Here, $\mu_{\mathrm{max}}$ is the source peak surface brightness above the background, and $\mathrm{mag}$ is the magnitude used to compute point-like probability, both given by SourceXtractor++ [^4] https://github.com/astrorama/SourceXtractorPlusPlus (Bertin et al. 2020; Kümmel et al. 2022) during the MER data reduction. The estimator $\mu_{\mathrm{max}}-\mathrm{mag}$ is related to the concentration of light at the peak versus the total magnitude; at a given magnitude, sources with smaller $\mu_{\mathrm{max}}-\mathrm{mag}$ are more point-like (Euclid Collaboration: Romelli et al. 2025). This parameter has also been used in Jauzac et al. (2012), Sharon et al. (2022), Estrada et al. (2023), Euclid Collaboration: Matamoro Zatarain et al. (2025), and Euclid Collaboration: Roster et al. (2025) as input for point/extended source classification.
@@
Figure 2 shows the distribution of visually identified quasars in the $\mu_{\mathrm{max}}-\mathrm{mag}$-$z_{\mathrm{vi}}$ plane. In general, $\mu_{\mathrm{max}}-\mathrm{mag}$ decreases as $z_{\mathrm{vi}}$ becomes higher. Most sources exhibit compact morphologies (low $\mu_{\mathrm{max}}-\mathrm{mag}$ values) at $z_{\mathrm{vi}}\gtrsim 0.8$, consistent with unresolved point-like sources. However, a distinct subset of sources at $z_{\mathrm{vi}}>2$ shows significantly higher $\mu_{\mathrm{max}}-\mathrm{mag}$ values, indicative of extended or low-concentration profiles that are atypical for high-redshift quasars. These sources are flagged as spurious redshifts and excluded from the final sample using a conservative empirical cut: $\mu_{\mathrm{max}}-\mathrm{mag}>-2$ at $z_{\mathrm{vi}}>2$, as indicated by the dashed red lines. This selection removes sources whose morphology and redshift are inconsistent with expectations for high-redshift quasars, improving the redshift accuracy of the final catalogue.
@@
Figure 2: Concentration parameter $\mu_{\mathrm{max}}-\mathrm{mag}$ as a function of the visual redshift $z_{\mathrm{vi}}$. The rejected region is shaded in pink, and is defined with $\mu_{\mathrm{max}}-\mathrm{mag}>-2$ at $z_{\mathrm{vi}}>2$, as indicated by the dashed red lines. Sources inside the rejected region are considered to have spurious redshifts and are marked with black crosses.
As an external consistency check on our visually confirmed redshifts, we crossmatched our quasar sample with Data Release 1 of the Dark Energy Spectroscopic Instrument (DESI DR1; DESI Collaboration: Abdul-Karim et al. 2025) using a radius of 1.0 and obtained 454 objects in common. We flag discrepant cases using the criterion $\left|z_{\mathrm{vi}}-z_{\mathrm{DESI}}\right|/(1+z_{\mathrm{DESI}})>0.15$, yielding 16 outliers. We then visually reinspect both the DESI and Euclid spectra for all 16 sources. In 11 cases, the DESI spectra support the same redshift solution as our Euclid-based identification, while the DESI catalogue redshift is inconsistent with the spectral features. In the remaining five cases, we revise our visual redshifts; the updated $z_{\mathrm{vi}}$ values are now consistent with $z_{\mathrm{DESI}}$ and are adopted in the released catalogue. The spectra of the 16 initially discrepant sources are shown in Appendix A.
The visually identified quasar sample contains 3468 sources (38 % of the entire sample) covering the redshift range of $0<z\lesssim 4.8$. In addition to the crossmatch to DESI DR1 above, we also crossmatch the full sample with the Million Quasar Catalogue (Milliquas v8; Flesch 2023) using a radius of 1.0 and find 710 sources in common. The union of the DESI DR1 matches (454 objects) and the Milliquas matches (710 objects) contain 782 sources, implying that 2686 quasars in our sample are new spectroscopic identifications.